OpenStudy (anonymous):

What are the solutions of x^2 = 8?

OpenStudy (abb0t):

\(\pm 2\sqrt{2}\)

OpenStudy (hunus):

\[\Large x^{2}=2^{3}\] \[\Large \sqrt{x^2} = x= \pm \sqrt{2^3}=\pm 2^{\frac{3}{2}}\]